OpenStudy (anonymous):

if the cost, C, for manufacturing x units of a certain product is given by C=x^2-5x+70, find the number of units manufactured at a cost of $8620

OpenStudy (richyw):

well \(C(x)=x^2-5x+70\) and you know the cost is $8620 so \[x^2-5x+70=8620\]now solve for x. Do you know how to do that?

OpenStudy (anonymous):

no, that is where I'm having the problem cause I don't understand algebra

OpenStudy (anonymous):



OpenStudy (anonymous):

quadratic formula, where a = 1, b= 5, c = -8550

OpenStudy (anonymous):

b = -5 sorry

OpenStudy (richyw):

alright well the first thing you usually want to do is write this in general form. To do this you subtract 8620 from both sides

OpenStudy (anonymous):

I have no idea what you are talking about

OpenStudy (richyw):

\[x^2-5x+70-8620=0\]\[x^2-5x-8850=0\]Now you want to find the roots to that. First you want to see if it can be factored, if it can then that's great and if not you need to use the quadratic formuala.

OpenStudy (richyw):

i'm getting there!

OpenStudy (anonymous):

ok

OpenStudy (richyw):

the quadratic formula finds the roots of quadratic polynomials. For a quadratic polynomial in general form, that is \(f(x)=ax^2+bx+c=0\), the quadratic formulas is \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]